# Topological equivalence and phase transition rate in holographic thermodynamics of regularized Maxwell theory

### v1

@@ I Introduction
where $G_{N}$ stands for the Newton’s constant and $l$ is the $D$ dimension AdS curvature radius. The thermodynamic volume is identified as the conjugate variable of pressure. By incorporating these new quantities, the extended form of the first law of thermodynamics and the corresponding Smarr relation for a charged AdS black hole can be expressed as follows
@@
In this new framework, the mass $M$ has no longer meaning of internal energy but reinterpreted as a chemical enthalpy Karatra2009 . In the extended phase space, the thermodynamics of these AdS black holes exhibit a wide range of phenomena, including Van der Waals-like phase transitions Kuma2012 ; Caca2013 ; Nam2018 , reentrant phase transitions Alkuma2013 ; Kurihe2021 ; Cuxuzhu2023 , black hole heat engine John2014 ; Bhaye2017 ; Zhali2019 ; Nam2021 , and Joule-Thompson expansions Okay2017 ; Molila2018 ; Nam2020 .
@@
$+$ $\left(M-\frac{\kappa A}{8\pi G_{N}}-\Omega J-\frac{\Phi}{l}Ql\right)\frac{d(l^{D-2}/G_{N})}{l^{D_{2}}/G_{N}}.$
@@
Models of non-linear electrodynamics (NLEs) have long been investigated as classical attempts to regularize the field of a point charge in Maxwell’s theory Boin1933 ; Soro2022 . The important idea was to modify the electromagnetic Lagrangian so that the point charge could exhibit a finite field strength and finite self-energy. Among these, perhaps the most renowned NLE theory is the Born-Infeld theory Boin1933 ; Boin1934 . This theory features finite self-energy for point charges, possesses electromagnetic duality, and approximates Maxwell’s electrodynamics in the weak field limit. It was later discovered that Born-Infeld-type Lagrangians also emerge in the low-energy regime of string theory Fratse1985 and in D-brane physics Leigh1989 . Recently, Einstein gravity coupled with NLEs has garnered significant scholarly attention due to its potential in addressing the singularity problem associated with black holes Bardeen1968 ; Beagar2000 ; Fanwa2016 .
In this paper, we explore the intriguing characteristics of the holographic thermodynamics of AdS black hole solutions within the regularized Maxwell theory (RegMax theory), a NLE theory first introduced in Taham2021 . This theory, which admits the Maxwell limit, offers the simplest regularization of the point charge field and its self-energy. Aside from Maxwell’s theory, it stands as a unique model of non-linear electrodynamics that provides radiative solutions in the Robinson-Trautman class. Remarkably, slowly rotating black holes or charged accelerated black holes are also naturally found in the RegMax theory Taham2021 ; Haku2023 . The spherical black hole solutions and their thermodynamic properties within RegMax theory were thoroughly investigated in Haku2023 . Their findings demonstrate that the characteristic parameter of RegMax theory significantly influences the thermodynamic behavior of the black holes.

### v2

@@ I Introduction
where $G_{N}$ stands for the Newton’s constant and $l$ is the $D$ dimension AdS curvature radius. The thermodynamic volume is identified as the conjugate variable of pressure. By incorporating these new quantities, the extended form of the first law of thermodynamics and the corresponding Smarr relation for a charged AdS black hole can be expressed as follows Kumate2017
@@
In this new framework, the mass $M$ has no longer meaning of internal energy but reinterpreted as a chemical enthalpy Karatra2009 . In the extended phase space, the thermodynamics of these AdS black holes exhibit a wide range of phenomena, including Van der Waals-like phase transitions Kuma2012 ; Caca2013 ; Nam2018 , reentrant phase transitions Alkuma2013 ; Kurihe2021 ; Cuxuzhu2023 , black hole heat engine John2014 ; Bhaye2017 ; Zhali2019 ; Nam2021 , and Joule-Thomson expansions Okay2017 ; Molila2018 ; Cihu2019 ; Nam2020 .
@@
$+$ $\left(M-\frac{\kappa A}{8\pi G_{N}}-\Omega J-\frac{\Phi}{l}Ql\right)\frac{d(l^{D-2}/G_{N})}{l^{D-2}/G_{N}}.$
@@
Models of non-linear electrodynamics (NLEs) have long been investigated as classical attempts to regularize the field of a point charge in Maxwell’s theory Boin1933 ; Soro2022 . The important idea was to modify the electromagnetic Lagrangian so that the point charge could exhibit a finite field strength and finite self-energy. Among these, perhaps the most renowned NLE theory is the Born-Infeld theory Boin1933 ; Boin1934 . This theory features finite self-energy for point charges, possesses electromagnetic duality, and approximates Maxwell’s electrodynamics in the weak field limit. It was later discovered that Born-Infeld-type Lagrangians also emerge in the low-energy regime of string theory Fratse1985 and in D-brane physics Leigh1989 . Recently, Einstein gravity coupled with NLEs has garnered significant scholarly attention due to its potential in addressing the singularity problem associated with black holes Bardeen1968 ; Beagar2000 ; Fanwa2016 . Interestingly, in the extended phase space, the nonlinear electrodynamic (NLE) AdS black holes also possesses thermodynamic behaviors such as Van der Waals-like phase transitions Guku2012 ; Kuli2018 ; Mabre2020 and Joule-Thomson expansions Bame2022 ; Krug20221 ; Krug20222 ; Krug2023 .
In this paper, we explore the intriguing characteristics of the holographic thermodynamics of AdS black hole solutions within the regularized Maxwell (RegMax) theory, a NLE theory first introduced in Taham2021 . This theory, which admits the Maxwell limit, offers the simplest regularization of the point charge field and its self-energy. Aside from Maxwell’s theory, it stands as a unique model of non-linear electrodynamics that provides radiative solutions in the Robinson-Trautman class. Notably, slowly rotating black holes or charged accelerated black holes within a given NLE framework, first constructed in the ModMax theory Baci2022 , are also naturally found in the RegMax theory Taham2021 ; Haku2023 . The spherical black hole solutions and their thermodynamic properties within RegMax theory were thoroughly investigated in Haku2023 . Their findings demonstrate that the characteristic parameter of RegMax theory significantly influences the thermodynamic behavior of the black holes.